# Elliptic asymptotic representation of the fifth Painlevé transcendents (corrected version)

### v8

@@ 1. Introduction
with $\theta_{0},$ $\theta_{1},$ $\theta_{\infty}\in\mathbb{C}$ defines nonlinear special functions, which are meromorphic on the universal covering space of $\mathbb{C}\setminus\{0\}$. A general solution is expressed by a convergent series in a spiral domain around $x=0$, and admits asymptotic representations as $x\to\infty$ along the real and the imaginary axes (cf. e.g. [37], [36]). Using the isomonodromy property and the WKB-analysis, for a generic case of (PV) Andreev and Kitaev [5] obtained families of solutions near $x=0$ and $x=\infty$ on the positive real axis, and connection formulas for these solutions. Along the imaginary axis asymptotic solutions and the associated monodromy data have been studied by [6], [38]. Furthermore Lisovyy et al. [29] gave a connection formula for the tau-function $\tau_{\mathrm{V}}(x)$ between $x=0$ and $x=i\infty$ and the ratios of multipliers of $\tau_{\mathrm{V}}(x)$ as $x\to 0,+\infty,i\infty$.
@@
which is equivalent to (PV) [19, Appendix C], [18], [5]; that is, $y(x)$ solves (PV) if and only if the monodromy data $M^{0},$ $M^{1},$ $S_{1},$ $S_{2}$ for (1.1) (defined in Section 2) remain invariant under a small change of $x.$ We apply the WKB-analysis to calculate the monodromy data as a direct monodromy problem, from which the elliptic expression of the main results follows. The main term expressed by the $\mathrm{sn}$-function with the modulus $A_{\phi}$ may be understood to depend on the phase shift only as a single integration constant parametrised by the monodromy data. The other integration constant is hidden in the error term, and is also deeply related to the correction function $B_{\phi}(t)=t(a_{\phi}(t)-A_{\phi}).$ The error term is expressed by an explicit asymptotic formula, whose leading term is written in terms of integrals of the $\mathrm{sn}$-function and the $\vartheta$-function, and contains the other integration constant [40, Theorems 2.2 and 2.3].
In the early version of [39] the Stokes graph is incorrect, and it affects the phase shifts of asymptotic solutions in [39, Theorems 2.1 and 2.2]. This paper contains the amended Stokes graph and the subsequent modifications including the corrected phase shifts in these theorems (Corrigendum of [39]).
@@
This paper is organised as follows. Section 2 describes our main results on the elliptic expression by the Jacobi $\mathrm{sn}$-function for $0<|\phi|<\pi/2$ with $\phi=\arg x$ (Theorem 2.1) and for $0<|\phi\mp\pi|<\pi/2$ (Theorem 2.2), and on the error term with the explicit leading term (Theorem 2.3 from [40]). Theorems 2.1 and 2.2 contain corrected phase shifts. Section 3 is devoted to basic facts necessary in proving the main results: parametrisation of $y(x)$ by the monodromy data; turning points and Stokes curves for the symmetric linear system (3.4) with $\lambda=e^{i\phi}(2\xi-1)$; and a WKB-solution in the canonical domain and local solutions around turning points. Section 4 is devoted to a direct monodromy problem. In the calculation of the monodromy we use the Stokes matrix and a connection matrix along a path passing through turning points. In other words, it is possible to replace the calculation of the monodromy matrices by that of suitable connection matrices. Our discussion follows the justification scheme of Kitaev [26], and the WKB analysis is carried out under the supposition (4.3). The monodromy matrices are expressed in term of integrals related to the characteristic roots and the WKB-solution. In Section 5, these integrals are represented by the elliptic integrals of the first and the second kind, and the $\vartheta$-function (Propositions 5.3, 5.5 and Corollary 5.4). In Section 6, considering an inverse monodromy problem by the use of the propositions in Section 5 with the justification scheme, we obtain the elliptic asymptotic representation of $y(x)$ as in Theorem 2.1 and an asymptotic expression of the correction function $B_{\phi}(t)$. For Theorem 2.2 the proof is sketched. Section 7 gives an alternative approach to $B_{\phi}(t)$, by using a system of equations equivalent to (PV) via its Lagrangian. This system is employed in the study of the error term [40]. Several properties of the Boutroux equations and the trajectory of its solution are important in our arguments, which are discussed in the final section.
In the preceding version the phase shift of Theorem 2.1 was corrected ($\Omega_{\mathbf{a}}/2$ is added). In this version, the following changes are made:
@@
(i) Theorem 2.2 is renewed to be more concretely. The statements of Theorems 2.3 and 2.4 are replaced with unconditional results cited from [40]. Consequently the proofs of these theorems of the early version are removed from Section 6.
@@
(ii) In the proof of Proposition 3.3 a minor revision is made. Estimates in Proposition 3.4 are improved and the proof are revised. Remarks 3.8 and 4.1 are added, which are related to Proposition 3.4.
@@ 2. Main results
as $\xi\to\infty$ through the sector $|\arg(x\xi)-\pi/2|<\pi$. Let $M^{0},M^{1},M^{\infty}\in SL_{2}(\mathbb{C})$ be the monodromy matrices given by the analytic continuations of $\Xi(\xi)$ along loops $l_{0},$ $l_{1}$, $l_{\infty}\in\pi_{1}(P^{1}(\mathbb{C})\setminus\{0,1,\infty\})$ as in Figure 2.1 defined for $\pi/2<\arg x<\pi/2$, which start from the point $p_{\mathrm{st}}$ satisfying $100<|p_{\mathrm{st}}|<\infty$ and $\arg(xp_{\mathrm{st}})=\pi/2$ and surround, respectively, $\xi=0$, $\xi=1$ and $\xi=\infty$ anticlockwise. It is easy to see $M^{\infty}M^{1}M^{0}=I.$
@@
These monodromy data $M^{0},$ $M^{1},$ $S_{1},$ $S_{2}$ are as given in [5, (2.3), (2.12)]. As will be explained in Section 3.1, if $M^{0},$ $M^{1}\not=\pm I$, solutions of (PV) are labelled and parametrised by entries of $M^{0},$ $M^{1},$ or of $S_{1},$ $S_{2}$.
For a given number $\phi\in\mathbb{R}$, the Boutroux equations
(2.3) $\mathrm{Re\,}e^{i\phi}\int_{\mathbf{a}}\sqrt{\frac{A_{\phi}-z^{2}}{1-z^{2}}}dz=\mathrm{Re\,}e^{i\phi}\int_{\mathbf{b}}\sqrt{\frac{A_{\phi}-z^{2}}{1-z^{2}}}dz=0$
admit a unique solution $A_{\phi}\in\mathbb{C}$ having the properties (cf. Proposition 8.17 and Figure 8.2, (a)):
@@
(iii) for $0\leq\phi\leq\pi/2,$ $\mathrm{Im\,}A_{\phi}\geq 0$, and, for $\pi/2\leq\phi\leq 0,$ $\mathrm{Im\,}A_{\phi}\leq 0$.
By (i) we fix $\mathrm{Re\,}A^{1/2}_{\phi}\geq 0,$ and then $(\overline{A_{\phi}})^{1/2}=\overline{A^{1/2}_{\phi}}.$ Here $\mathbf{a}$ and $\mathbf{b}$ denote basic cycles as in Figure 2.2 on the elliptic curve $\Pi^{*}=\Pi^{*}_{+}\cup\Pi^{*}_{-}$ given by $w(A_{\phi},z)=\sqrt{(1-z^{2})(A_{\phi}-z^{2})}$ such that $\Pi^{*}_{+}$ and $\Pi^{*}_{-}$ are glued along the cuts $1,-A_{\phi}^{1/2}]$ and $A_{\phi}^{1/2},1]$. The branches of these algebraic functions are fixed in such a way that $z^{-2}\sqrt{(A_{\phi}-z^{2})(1-z^{2})}\to-1$ and $\sqrt{(A_{\phi}-z^{2})/(1-z^{2})}\to 1$ as $z\to\infty$ on the upper sheet $\Pi^{*}_{+}.$ Write
@@
$A_{\phi}^{1/2}11$$A_{\phi}^{1/2}$$\mathbf{a}$$\mathbf{b}$$\Pi^{*}_{+}$
Figure $2.$ Cycles $\mathbf{a},$ $\mathbf{b}$ on $\Pi^{*}$
@@ 2.1. Elliptic representation
Suppose that $0<|\phi|<\pi/2.$ Let $y(x)$ be the solution of $(\mathrm{P}_{\mathrm{V}})$ corresponding to $M^{0}=(m_{ij}^{0}),$ $M^{1}=(m_{ij}^{1})$ such that $m_{11}^{0}m_{11}^{1}m_{21}^{0}m_{12}^{1}\not=0.$ Then
@@
$\delta>0$ being some positive number, $\kappa_{0}>0$ a given number, $\delta_{0}>0$ a given small number, and $t_{\infty}=t_{\infty}(\kappa_{0},\delta_{0})$ a large number depending on $(\kappa_{0},\delta_{0}).$ Furthermore $x_{0}$ is such that $x_{0}\in S(\phi,t_{\infty},\kappa_{0},\delta_{0})$ and that
@@
The conclusion of Theorem 2.1 for $\pi/2<\phi<0$ (respectively, $0<\phi<\pi/2$) remains valid under the weaker condition $m^{0}_{11}m^{0}_{21}m^{1}_{12}\not=0$ (respectively, $m^{1}_{11}m^{0}_{21}m^{1}_{12}\not=0$), since the passage of the limit $m^{1}_{11}\to 0$ (respectively, $m^{0}_{11}\to 0$) is possible in (4.4). The complementary case, say $m^{1}_{11}m^{0}_{21}m^{1}_{12}=0$ for $0<\phi<\pi/2$ corresponds to truncated or triply-truncated solutions [5, Corollaries 5.2 and 5.3], [4, §5], [2, Theorem 1.1], [38, Theorem 2.21] (see also Remark 3.6). Similar relaxing of a condition may be made in Theorem 2.2 below.
@@
Among the singular values $y=0,$ 11, $\infty$ of (PV), the neighbourhoods of only $1$-points are excluded from the cheese-like domain of the elliptic expression of $y(x)$.
@@
Our calculation of Section 4 leads to $\delta=1/4-\varepsilon$ for any $0<\varepsilon<1/4$, whose numerical value is caused by a technical reason. In fact we have $\delta=1$ by Theorem 2.3, [40, Theorem 2.1].
The solution $y(x)$ given in Theorem 2.1 is labelled by $M^{0,\,1}=(m^{0,\,1}_{ij})$. Denote this by $y(x)=y(M^{0},M^{1};x).$ Note that $\Omega_{\mathbf{a}}$ and $\Omega_{\mathbf{b}}$ are determined by $A_{\phi}$, which does not depend on $M^{0,\,1}$, and satisfy $\Omega_{\mathbf{a},\mathbf{b}}(\phi+\pi)=\Omega_{\mathbf{a},\mathbf{b}}(\phi)$.
@@
Let $0<|\phi-\pi|<\pi/2.$ Write $\breve{M}^{0}=S^{-1}_{2}M^{0}S_{2},$ $\breve{M}^{1}=S^{-1}_{2}M^{1}S_{2}$ with $\breve{M}^{0}=(\breve{m}^{0}_{ij}),$ $\breve{M}^{1}=(\breve{m}^{1}_{ij}).$ If $\breve{m}^{0}_{12}\breve{m}^{1}_{21}\breve{m}^{0}_{22}\breve{m}^{1}_{22}\not=0,$ then $y(x)=y(M^{0},M^{1};x)$ admits an elliptic representation of the form
@@
$\breve{x}_{0}$ $\equiv\frac{-1}{\pi i}\Bigl(\Omega_{\mathbf{b}}\log(m^{0}_{21}m^{1}_{12})+\Omega_{\mathbf{a}}\log\breve{\mathfrak{m}}_{\phi}\Bigr)-\Bigl(\frac{\Omega_{\mathbf{a}}}{2}+\Omega_{\mathbf{b}}\Bigr)(\theta_{\infty}+1)-\frac{\Omega_{\mathbf{a}}}{2}$
@@
with $\breve{\mathfrak{m}}_{\phi,\theta_{\infty}}=e^{\pi i\theta_{\infty}/2}(\breve{m}^{0}_{22})^{-1}$ if $\pi/2<\phi<\pi,$ and $e^{-\pi i\theta_{\infty}/2}\breve{m}^{1}_{22}$ if $\pi<\phi<3\pi/2,$ and $\breve{\mathfrak{m}}=e^{-\pi i\theta_{\infty}/2}\breve{\mathfrak{m}}_{\phi,\theta_{\infty}}.$
@@
Under a weaker condition $\breve{m}^{0}_{12}\breve{m}^{1}_{21}\breve{m}^{0}_{22}\not=0$ (respectively, $\breve{m}^{0}_{12}\breve{m}^{1}_{21}\breve{m}^{1}_{22}\not=0$), the asymptotic formula of the theorem is also valid in the sector $\pi/2<\phi<\pi$ (respectively, $\pi<\phi<3\pi/2$).
Remark 2.5.
Let $0<|\phi-2p\pi|<\pi/2$ or $|\phi-2p\pi-\pi|<\pi/2$ with $p\in\mathbb{Z}\setminus\{0\}$. Set
@@
(2) for $0<|\phi-2p\pi-\pi|<\pi/2,$ $(m^{0}_{12}m^{1}_{21})^{-1}\mapsto((\breve{m}^{0}_{p})_{12}(\breve{m}^{1}_{p})_{21})^{-1},$ $\breve{\mathfrak{m}}_{\phi}\mapsto\breve{\mathfrak{m}}^{p}_{\phi}=\breve{\mathfrak{m}}_{\phi}|_{\breve{m}^{0,1}_{22}\mapsto(\breve{m}^{0,1}_{p})_{22}}$ in Theorem 2.2.
The facts in the remark above suggest the nonlinear monodromy and Stokes structure of (PV), which will be discussed elsewhere.

### v9

@@ 1. Introduction
with $\theta_{0},$ $\theta_{1},$ $\theta_{\infty}\in\mathbb{C}$ defines nonlinear special functions, which are meromorphic on the universal covering space of $\mathbb{C}\setminus\{0\}$. A general solution is expressed by a convergent series in a spiral domain around $x=0$, and admits asymptotic representations as $x\to\infty$ along the real and the imaginary axes (cf. e.g. [37], [36]). Using the isomonodromy property and the WKB analysis, for a generic case of (PV) Andreev and Kitaev [5] obtained families of solutions near $x=0$ and $x=\infty$ on the positive real axis, and connection formulas for these solutions. Along the imaginary axis asymptotic solutions and the associated monodromy data have been studied by [6], [38]. Furthermore Lisovyy et al. [29] gave a connection formula for the tau-function $\tau_{\mathrm{V}}(x)$ between $x=0$ and $x=i\infty$ and the ratios of multipliers of $\tau_{\mathrm{V}}(x)$ as $x\to 0,+\infty,i\infty$.
@@
which is equivalent to (PV) [19, Appendix C], [18], [5]; that is, $y(x)$ solves (PV) if and only if the monodromy data $M^{0},$ $M^{1},$ $S_{1},$ $S_{2}$ for (1.1) (defined in Section 2) remain invariant under a small change of $x.$ We apply the WKB analysis to calculate the monodromy data, which is used in deriving the elliptic expression of the main results. The main term expressed by the $\mathrm{sn}$-function may be understood to depend on the phase shift only as a single integration constant parametrised by the monodromy data. The other integration constant is hidden in the error term, and is also deeply related to the correction function $B_{\phi}(t)=t(a_{\phi}(t)-A_{\phi})$ with $\phi=\arg x.$ The error term admits an explicit asymptotic formula containing the other integration constant [40, Theorems 2.2 and 2.3].
@@
This paper is organised as follows. Section 2 describes our main results on the expression of a general solution by the Jacobi $\mathrm{sn}$-function for $0<|\phi|<\pi/2$ (Theorem 2.1) and for $0<|\phi-\pi|<\pi/2$ (Theorem 2.2). Section 3 provides basic facts necessary in proving the main results: parametrisation of $y(x)$ by the monodromy data; turning points and Stokes curves for the symmetric linear system (3.4) with $\lambda=e^{i\phi}(2\xi-1)$; and a WKB-solution in the canonical domain and local solutions around turning points. Section 4 is devoted to a direct monodromy problem for system (3.4). The monodromy matrices are obtained from connection matrices along suitably chosen paths, which are calculated by WKB analysis under the supposition (4.3). The monodromy matrices are expressed in term of integrals related to the characteristic roots and the WKB-solutions. In Section 5, these integrals are represented by elliptic integrals and the $\vartheta$-function (Propositions 5.3, 5.5 and Corollary 5.4). In Section 6, considering an inverse monodromy problem by the use of the propositions in Section 5 and following the justification scheme by Kitaev [26], we arrive at the elliptic asymptotic representation of $y(x)$ as in Theorem 2.1 and an asymptotic expression of the correction function $B_{\phi}(t)$. For Theorem 2.2 the proof is sketched. Section 7 gives an alternative approach to $B_{\phi}(t)$, by using a system equivalent to (PV) via its Lagrangian. This system is employed in the study of the error term [40]. Several properties of the Boutroux equations and its solution are important in our arguments, which are discussed in the final section.
In the early version of [39] the Stokes graph is incorrect, and it affects the phase shifts of asymptotic solutions in [39, Theorems 2.1 and 2.2]. The amended Stokes graph and the subsequent modifications including the corrected phase shifts are contained in the second version (Corrigendum of [39]). In the preceding version, the further correction of the phase shift in Theorem 2.1 ($\Omega_{\mathbf{a}}/2$ is added), and the following changes were made:
@@
(i) Theorem 2.2 is renewed. Theorems 2.3 and 2.4 are replaced with unconditional results cited from [40], and the early proofs are removed from Section 6.
@@
(ii) In the proof of Proposition 3.4 a minor revision is made. Estimates in Proposition 3.5 are improved. Remarks 3.7 and 4.1 are added, which are related to Proposition 3.5.
@@
This version contains several minor revisions for readability, e.g. Theorem 2.1.
@@ 2. Main results
as $\xi\to\infty$ through the sector $|\arg(x\xi)-\pi/2|<\pi$. Let $M^{0},M^{1},M^{\infty}\in SL_{2}(\mathbb{C})$ be the monodromy matrices defined by $\Xi^{(l_{\nu})}(\xi)=\Xi(\xi)M^{\nu}$ for $\nu=0,1,\infty.$ Here $\Xi^{(l_{\nu})}(\xi)$, $\nu=0,1,\infty$ denote the analytic continuations of $\Xi(\xi)$ along the respective loops $l_{0},$ $l_{1}$, $l_{\infty}\in\pi_{1}(P^{1}(\mathbb{C})\setminus\{0,1,\infty\})$ as in Figure 2.1 defined for $\pi/2<\arg x<\pi/2$, which start from the point $p_{\mathrm{st}}$ satisfying $100<|p_{\mathrm{st}}|<\infty$ and $\arg(xp_{\mathrm{st}})=\pi/2$ and surround, respectively, $\xi=0$, $\xi=1$ and $\xi=\infty$ anticlockwise. It is easy to see $M^{\infty}M^{1}M^{0}=I.$
@@
The monodromy data $M^{0}=(m^{0}_{ij})$ and $M^{1}=(m^{1}_{ij})$ satisfy (3.2) [5]. Set
$\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}:=\{(M^{0},M^{1})\in SL_{2}(\mathbb{C})^{2}\,|\,$ $m^{1}_{11}m^{0}_{11}+m^{1}_{12}m^{0}_{21}=e^{-\pi i\theta_{\infty}},\,$
$\mathrm{tr}\,M^{0}=2\cos\pi\theta_{0},\,\mathrm{tr}\,M^{1}=2\cos\pi\theta_{1}\}/\sim,$
where $(M^{0},M^{1})\sim(\tilde{M}^{0},\tilde{M}^{1})$ if $(M^{0},M^{1})=c^{\sigma_{3}}(\tilde{M}^{0},\tilde{M}^{1})c^{-\sigma_{3}}$ for some $c\in\mathbb{C}\setminus\{0\}.$ Then $\mathrm{dim}_{\mathbb{C}}\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}=2.$ Let us call $\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}$ the monodromy manifold. As shown in Proposition 3.1, for every $(M^{0},M^{1})\in\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}$ such that $M^{0},$ $M^{1}\not=\pm I$, there exists uniquely a solution $y(x)$ of (PV) corresponding to $(M^{0},M^{1})$, and then $y(x)$ is labelled with $(M^{0},M^{1})$.
For $w(A,z):=\sqrt{(1-z^{2})(A-z^{2})}$ consider the Riemann surface $\Pi_{A}=\Pi_{+}\cup\Pi_{-}$ such that $\Pi_{+}$ and $\Pi_{-}$ are glued along the cuts $1,-A^{1/2}]$ and $A^{1/2},1]$ with $\mathrm{Re\,}A^{1/2}\geq 0$. Let the branches be such that $z^{-2}\sqrt{(1-z^{2})(A-z^{2})}\to-1$ and $\sqrt{(A-z^{2})/(1-z^{2})}\to 1$ as $z\to\infty$ on the upper sheet $\Pi_{+}.$ [^1] Then $A^{-1/2}\sqrt{(1-z^{2})(A-z^{2})}\to 1$ and $A^{-1/2}\sqrt{(A-z^{2})/(1-z^{2})}\to 1$ as $z\to 0$ on $\Pi_{+}.$ Let $\mathbf{a}$ and $\mathbf{b}$ denote basic cycles on $\Pi_{A}$ as drawn in Figure 2.2 when $A=A_{\phi}.$ For a given number $\phi\in\mathbb{R}$, the Boutroux equations
(2.3) $\mathrm{Re\,}e^{i\phi}\int_{\mathbf{a}}\sqrt{\frac{A-z^{2}}{1-z^{2}}}dz=\mathrm{Re\,}e^{i\phi}\int_{\mathbf{b}}\sqrt{\frac{A-z^{2}}{1-z^{2}}}dz=0$
admit a unique solution $A=A_{\phi}\in\mathbb{C}$ having the properties:
@@
(iii) for $0\leq\phi\leq\pi/2,$ $\mathrm{Im\,}A_{\phi}\geq 0$, and, for $\pi/2\leq\phi\leq 0,$ $\mathrm{Im\,}A_{\phi}\leq 0$
(cf. Proposition 8.17 and Figure 8.2, (a)). By (i) we fix $\mathrm{Re\,}A^{1/2}_{\phi}\geq 0,$ and then $(\overline{A_{\phi}})^{1/2}=\overline{A^{1/2}_{\phi}}.$ In what follows let $\mathbf{a}$ and $\mathbf{b}$ be on $\Pi_{A_{\phi}}=\Pi_{+}\cup\Pi_{-}$ as in Figure 2.2. Write
@@
$A_{\phi}^{1/2}11$$A_{\phi}^{1/2}$$\mathbf{a}$$\mathbf{b}$$\Pi_{+}$
Figure 2.2. Cycles $\mathbf{a},$ $\mathbf{b}$ on $\Pi_{A_{\phi}}$
@@ 2.1. Elliptic representation
Let $M^{0}=(m_{ij}^{0})$ and $M^{1}=(m_{ij}^{1})$ be such that $m_{11}^{0}m_{21}^{0}m_{12}^{1}\not=0$ (respectively, $m_{11}^{1}m_{21}^{0}m_{12}^{1}\not=0$). Then, for $\pi/2<\phi<0$ (respectively, $0<\phi<\pi/2$), a solution $y(x)$ labelled with $(M^{0},M^{1})$ admits an elliptic representation of the form
@@
$\delta>0$ being some positive number, $\kappa_{0}>0$ a given number, $\delta_{0}>0$ a given small number, and $t_{\infty}=t_{\infty}(\kappa_{0},\delta_{0})$ a large number depending on $(\kappa_{0},\delta_{0}).$ Furthermore $x_{0}$ is such that
@@
In Theorem 2.1 the complementary cases, say $m^{1}_{11}m^{0}_{21}m^{1}_{12}=0$ for $0<\phi<\pi/2$ correspond to truncated or triply-truncated solutions [5, Corollaries 5.2 and 5.3], [4, §5], [2, Theorem 1.1], [38, Theorem 2.21] (see also Remark 3.5). In Theorem 2.2 below similar truncated solutions exist in the complementary cases.
@@
The pair $(\mathfrak{m}_{\phi},m^{0}_{21}m^{1}_{12})$ uniquely determines $(M^{0},M^{1})\in\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}$.
@@
The point $x=\rho$ is in $\mathcal{P}_{0}$ if and only if $y(\rho)=1.$
Remark 2.4.
Our calculation of Section 4 leads to $\delta=1/4-\varepsilon$ for any $0<\varepsilon<1/4$, whose numerical value is caused by a technical reason. In fact we have $\delta=1$ in Theorem 2.3, [40, Theorem 2.1].
The solution in Theorem 2.1 labelled with $(M^{0},M^{1})$ may be written as $y(x,M^{0},M^{1})$. Let us denote by the same symbol $y(x,M^{0},M^{1})$ the analytic continuation to the sectors $0<|\phi-\pi|<\pi/2$. Note that $\Omega_{\mathbf{a}}$ and $\Omega_{\mathbf{b}}$ are determined by $A_{\phi}$, which does not depend on $M^{0},M^{1}$, and satisfy $\Omega_{\mathbf{a},\mathbf{b}}(\phi+\pi)=\Omega_{\mathbf{a},\mathbf{b}}(\phi)$.
@@
Write $\breve{M}^{0}=(\breve{m}^{0}_{ij})=S^{-1}_{2}M^{0}S_{2}$ and $\breve{M}^{1}=(\breve{m}^{1}_{ij})=S^{-1}_{2}M^{1}S_{2}$. Suppose that $\breve{m}^{0}_{22}\breve{m}^{0}_{12}\breve{m}^{1}_{21}\not=0$ (respectively, $\breve{m}^{1}_{22}\breve{m}^{0}_{12}\breve{m}^{1}_{21}\not=0$). Then, for $\pi/2<\phi<\pi$ (respectively, $\pi<\phi<3\pi/2$), $y(x)=y(x,M^{0},M^{1})$ admits an elliptic representation of the form
@@
$\breve{x}_{0}$ $\equiv\frac{-1}{\pi i}\Bigl(\Omega_{\mathbf{b}}\log(\breve{m}^{0}_{12}\breve{m}^{1}_{21})+\Omega_{\mathbf{a}}\log\breve{\mathfrak{m}}_{\phi}\Bigr)-\Bigl(\frac{\Omega_{\mathbf{a}}}{2}+\Omega_{\mathbf{b}}\Bigr)(\theta_{\infty}+1)-\frac{\Omega_{\mathbf{a}}}{2}$
@@
with $\breve{\mathfrak{m}}_{\phi,\theta_{\infty}}=e^{\pi i\theta_{\infty}/2}(\breve{m}^{0}_{22})^{-1}$ if $\pi/2<\phi<\pi,$ and $e^{-\pi i\theta_{\infty}/2}\breve{m}^{1}_{22}$ if $\pi<\phi<3\pi/2,$ and $\breve{\mathfrak{m}}_{\phi}=e^{-\pi i\theta_{\infty}/2}\breve{\mathfrak{m}}_{\phi,\theta_{\infty}}.$
@@
Let $0<|\phi-2p\pi|<\pi/2$ or $0<|\phi-2p\pi-\pi|<\pi/2$ with $p\in\mathbb{Z}\setminus\{0\}$. Set
@@
(2) for $0<|\phi-2p\pi-\pi|<\pi/2,$ $(\breve{m}^{0}_{12}\breve{m}^{1}_{21})^{-1}\mapsto((\breve{m}^{0}_{p})_{12}(\breve{m}^{1}_{p})_{21})^{-1},$ $\breve{\mathfrak{m}}_{\phi}\mapsto\breve{\mathfrak{m}}^{p}_{\phi}=\breve{\mathfrak{m}}_{\phi}|_{\breve{m}^{0,1}_{22}\mapsto(\breve{m}^{0,1}_{p})_{22}}$ in Theorem 2.2.